# HumanoidBench: Simulated Humanoid Benchmark for Whole-Body Locomotion and Manipulation

### v1

@@ Abstract
Humanoid robots hold great promise in assisting humans in diverse environments and tasks, due to their flexibility and adaptability leveraging human-like morphology. However, research in humanoid robots is often bottlenecked by the costly and fragile hardware setups. To accelerate algorithmic research in humanoid robots, we present a high-dimensional, simulated robot learning benchmark, HumanoidBench, featuring a humanoid robot equipped with dexterous hands and a variety of challenging whole-body manipulation and locomotion tasks. Our findings reveal that state-of-the-art reinforcement learning algorithms struggle with most tasks, whereas a hierarchical learning baseline achieves superior performance when supported by robust low-level policies, such as walking or reaching. With HumanoidBench, we provide the robotics community with a platform to identify the challenges arising when solving diverse tasks with humanoid robots, facilitating prompt verification of algorithms and ideas. The open-source code is available at https://sferrazza.cc/humanoidbench_site.
@@
TABLE I: Comparison of simulated robotic manipulation benchmarks. Our humanoid robot benchmark tests a variety of complex, long-horizon task with a large action space.
@@ I Introduction
To accelerate the progress of research for humanoid robots, we present the first-of-its-kind humanoid robot benchmark, HumanoidBench, with a diverse set of locomotion and manipulation tasks. Our simulated humanoid benchmark demonstrates a variety of challenges in addressing learning for autonomous humanoid robots, such as the intricate control of robots with complex dynamics, sophisticated coordination among various body parts, and addressing long-horizon complex tasks, while providing an accessible, fast, safe, and inexpensive testbed to robot learning researchers.
HumanoidBench contributes (1) a simulation environment comprising a humanoid robot with two dexterous hands, as illustrated in LABEL:fig:teaser; (2) a variety of tasks, spanning locomotion, manipulation, and whole-body control, incorporating humans’ everyday tasks; (3) a standardized benchmark to evaluate the progress of the community on high-dimensional humanoid robot learning and control; and finally, (4) benchmarking results of the state-of-the-art reinforcement learning (RL) algorithms and a hierarchical RL approach.
The simulation environment of HumanoidBench uses the MuJoCo [52] physics engine. For the simulated humanoid robot, we mainly opt for a Unitree H1 humanoid robot [^1] https://www.unitree.com/h1, which is relatively affordable and offers accurate simulation models [58], with two dexterous Shadow Hands [^2] https://www.shadowrobot.com/dexterous-hand-series/ attached to its arms. Our environment can easily incorporate any humanoid robots and end effectors; thus, we provide other models, including Agility Robotics Digit [^3] https://agilityrobotics.com/robots and the Robotiq 2F-85 gripper.
@@ II Related Work
Complex manipulation tasks, such as block stacking [11], kitchen tasks [14], and table-top manipulation [22, 34, 29], have been introduced but are still limited to a combination of pushing, picking, and placing. On the other hand, the IKEA furniture assembly environment [27], BEHAVIOR [48, 28], and Habitat [49] present diverse long-horizon (mobile) manipulation tasks, with their main focus being on high-level planning by abstracting complex low-level control problems, while FurnitureBench [19] introduces a simulated benchmark for complex long-horizon furniture assembly tasks with sophisticated low-level control. However, most of these benchmarks use a single-arm manipulation setup with either a parallel gripper or a dexterous hand [8, 42], limiting the types of object interactions and not addressing the challenges of coordinating multiple parts of a body [26], e.g., multiple fingers, arms, and legs.
Robosuite [62] includes a handful of bimanual manipulation tasks, while more recently [9] and [59] have introduced additional benchmarks that require coordinating two floating robot hands, i.e., not attached to any arm base. While bimanual manipulation is one of the key objectives of humanoid robots, most benchmarks in humanoid research have so far focused on the locomotion challenges of such platforms [7, 24, 39, 3]. In this regard, such simulations have accelerated research on control algorithms [5, 40, 41, 35], ultimately leading to achieve robust humanoid locomotion in the real world [1, 43].
@@
Figure 0: Example egocentric visual (top-left) and whole-body tactile (right) observations when the humanoid interacts with a package in the truck environment. In the right figure, the two cameras on the robot head are highlighted in green, while continuous tactile pressure readings are indicated in shades of red and yellow. Note that here, for ease of visualization, we are not showing shear forces and tactile readings on the back of the robot, which are also implemented in our environment.
@@
Figure 1: HumanoidBench manipulation task suite. We devise $15$ benchmarking whole-body manipulation tasks that cover a wide variety of interactions and difficulties. This figure illustrates an initial state (left) and target state (right) for each task.
@@
Figure 2: HumanoidBench locomotion task suite. We devise $12$ benchmarking locomotion tasks that cover a wide variety of interactions and difficulties. This figure illustrates an initial state (left) and target state (right) for each task, except for walk.
@@ III Simulated Humanoid Robot Environment
In this section, we describe our simulated environment and discuss relevant design choices for the simulated humanoid robot. As illustrated in Figure , we use the Unitree H1 humanoid robotFootnote 1 with two dexterous Shadow HandsFootnote 2 as the primary robotic agent of our benchmark. We simulate this humanoid robot using MuJoCo [52] adapting the Unitree H1 model provided by Unitree [^4] https://github.com/unitreerobotics/unitree_ros and the dexterous Shadow Hand models available through MuJoCo Menagerie [58].
Humanoid Body. We implement Unitree H1Footnote 1 and Agility Robotics DigitFootnote 3, which are well-known humanoid robots with their model files freely available [58]. Unitree H1 is primarily used in our benchmark as we observed faster learning, which we ascribe to a simpler mechanical design compared to Agility Robotics Digit, which features passive joints actuated through a four-bar linkage.
Dexterous Hands. We use two dexterous Shadow HandsFootnote 2, which also have model files freely available [58], and have shown impressive manipulation capabilities both in simulation [59] and the real world [2]. To make the simulated robot have more human-like morphology, we remove the cumbersome forearms of the dexterous Shadow Hands in HumanoidBench. While this is not currently a realistic model, we anticipate the trend in the industry towards developing slimmer, human-like hands (e.g. Tesla Optimus, Figure 01) so that our design choice aligns better with next-generation humanoid robots. In addition, we also provide models for the Robotiq 2F-85 parallel-jaw gripper and the hands available in the Unitree collectionFootnote 4.
@@
  • Proprioceptive robot state (i.e. joint angles and velocities) and task-relevant environment observations (i.e. object poses and velocities).
@@
  • Whole-body tactile sensing using the MuJoCo tactile grid sensor (see Figure ). We design tactile sensing at the hands with high resolution and in other body parts with low resolution, similar to humans, with a total of $448$ taxels spread over the entire body, each providing three-dimensional contact force readings. Similar distributed force readings have been captured on real-world systems both on humanoid bodies [36] and end-effectors [46]. The implementation of such spatially distributed contact sensing required non-trivial mesh adaptations and refinements, which we detail in the appendix.
Although other sensory inputs are available from the environment, to investigate challenges in whole-body control of humanoid robots, we first focus on the state-based environment setup, where proprioceptive robot states and object states are used for the agent’s input in HumanoidBench. We leave extending our environment to benchmarking multimodal perception capabilities [57, 47] of humanoid robots as future work. Note that in our state-based environment, we essentially maintain the robot observations the same across tasks to minimize domain knowledge, in contrast to tailoring it to the specific tasks [51].
@@
Figure 3: Learning curves of RL algorithms (locomotion). The curves are averaged over three random seeds and the shaded regions represent the standard deviation. The dashed lines qualitatively indicate task success. We run PPO on the walk task but it is not visible in the plot since it only achieves very low returns.
@@ IV HumanoidBench
To this end, we present HumanoidBench, a humanoid benchmark for robot learning and control, which features a high-dimensional action space (up to $61$ different actuator) and enables research in complex whole-body coordination.
We benchmark $27$ tasks, consisting of $12$ locomotion tasks and $15$ distinct manipulation tasks, as illustrated in Figure 2 and Figure 1. A set of locomotion tasks aim to provide interesting but simpler humanoid control scenarios, bypassing intricate dexterous hand control. On the other hand, whole-body manipulation tasks render a comprehensive evaluation of the state-of-the-art algorithms on challenging tasks with unique challenges that require coordination across the entire robot body, ranging from toy examples (e.g. pushing a box on a table) to practical applications (e.g. truck unloading, shelf rearrangement).
Below we briefly describe the tasks that are part of the benchmark. Further details about each of the tasks, including task initialization and reward functions, are provided in the appendix.
@@ IV-B Whole-Body Manipulation Tasks
  • cabinet: Open four different types of cabinet doors (e.g. hinge doors, sliding door, drawer).
@@
  • powerlift: Lifting a barbell shaped object of a designated mass.
@@ V Benchmarking Results
To identify the challenges in learning with humanoid robots, we benchmark reinforcement learning (RL) algorithms on HumanoidBench, which promises for robots to learn from their own experience. Remarkably, this class of algorithms requires limited domain expertise and does not necessarily rely on expert demonstrations, which are not only expensive but also challenging to collect for humanoid robots.
@@ V-A Baselines
We evaluate all tasks in our benchmark with four reinforcement learning methods (DreamerV3, TD-MPC2, SAC, PPO). Please refer to appendix, Appendix C, for implementation details.
@@ V-B Results
We report benchmarking results in Figure 3 and Figure 4, where we ran each of the algorithms for approximately 48 hours, resulting in the visible differences in environment steps (e.g. $2$M steps for TD-MPC2, $10$M steps for DreamerV3). We only run PPO on a subset of tasks (walk, kitchen, door, package), given its inferior performance without massive parallelization. Each of the environments is evaluated with a combination of dense rewards and sparse subtask completion rewards, and for each of these we provide qualitative measures of task success (see dashed lines in Figure 3 and Figure 4).
@@
This poor performance is mainly attributed to the high-dimensionality of the state and action spaces of our humanoid robot agent with dexterous hands. Although the hands of the humanoid robot are barely used for most locomotion tasks, the RL algorithms fail to ignore these information, which makes policy learning challenging. In addition, these high-dimensional state and action spaces result in a much larger exploration space, which makes exploration slow or infeasible with simple maximum entropy approaches. This implies the need for incorporating behavioral priors or common sense knowledge about the world that can ease the exploration problem, when it comes to learning more complex agents, like humanoid robots. We investigate this problem further in Section V-C.
This problem becomes even more severe in manipulation tasks, resulting in particularly low rewards in all manipulation tasks. Before learning any manipulation skills, an agent must learn locomotion skills to balance and move towards an object or the world to interact. All the policies barely learn to stabilize using the dense reward, but struggle to learn any complex manipulation skills.
@@
V-C With Hands vs. Without Hands
As discussed in Section V-B, controlling humanoid robots with dexterous hands is challenging due to their high degrees of freedom and complex dynamics. Thus, we investigate the difficulty of RL training with a large action space (i.e. additional $42$-DoFs with two dexterous Shadow Hands) on walk that does not necessarily require to control dexterous hands. The results in Figure 5 show that the presence of hands, with their additional joints and actuators, leads to a large decrease in performance compared to training the same task without the dexterous hands.
We perform an additional experiment to further confirm whether such difficulties stem from the dimensionality of the action space. In particular, we benchmark our full robot model, but fix the actuation of the hands, which we set to zero. In this way, the action dimensionality is reduced from $61$ in the original model to $19$, but the additional observations and masses induced by the presence of the hands are retained. On this experiment, the algorithms only slightly decrease their performance (see Blocked in the figure), confirming that most of the performance drop is indeed due to the increased action dimensionality. We repeat the same experiment, with similar results, by also removing the hand joints from the observation space (see Reduced in the figure). We observe similar trends in the more complex push task, which however, also presents substantially different dynamics in the task approach (e.g. pushing with and without hands).
@@
Figure 6: Our hierarchical architecture, where a robust low-level reaching policy (trained using PPO) follows targets provided by a high-level planning policy.
@@
As shown above, flat, end-to-end RL approaches fail to learn most of the tasks in HumanoidBench. The results above show that even state-of-the-art RL algorithms struggle on the most complex tasks in the benchmark. Many of such tasks require long-horizon planning and necessitate acquiring a diverse set of skills (e.g., balancing, walking, reaching, etc.) to successfully achieve the desired objective. We argue that these issues can be mitigated by introducing additional structure into the learning problem. In particular, we explore a hierarchical learning paradigm, where one or multiple low-level skill policies are provided to a high-level planning policy that sends setpoints to lower-level policies, see Figure 6. In practice, such setpoints comprise the action space of the high-level policy. This framework is very general, and there are no constraints on how to obtain both low-level and high-level policies. However, here we focus on training both of these through reinforcement learning.
@@
We report results of such a hierarchical approach on two different tasks, namely, the push and package tasks. As a low-level skill, push uses a one-hand reaching policy, which allows the robot to reach a 3D point in space with its left hand, while package uses a two-hand reaching policy, where both hands are commanded to reach different 3D targets.
@@
Ideally, the low-level skill policies are pretrained blocks that can be reused across tasks. In particular, our hierarchical architecture considers the low-level policy as a frozen block that does not improve during training of the high-level policy. As such, the low-level reaching policies in our push and package tasks need to be very robust to cope with the continually shifting reaching targets that the high-level policy sets during exploration. However, the results in the previous section show how even a one-hand reaching task is hard to learn for the algorithms benchmarked in this work.
@@
On a separate note, while our experiments above confirm that the on-policy PPO exhibits poor sample efficiency compared to the other off-policy algorithms, it is worth noting that PPO has achieved significant success in robotic locomotion by exploiting large-scale parallelization of environments on GPUs [31]. We exploit hardware acceleration using the recently released MuJoCo MJX [^5] https://mujoco.readthedocs.io/en/stable/mjx.html, which enables training PPO on thousands of parallel environments. We employ a simplified H1 model that only considers collisions between feet and ground in our MJX experiments, as in our experience the advantages stemming from parallelization are largely reduced when considering all numerous humanoid geometries (hindering training of the more complex benchmark tasks via MJX). We also remove the hands from the model to further increase training efficiency. We modify the reaching task by resetting the target during each episode whenever this has been reached. We also apply force perturbations at each of the links during training. We train the one-hand reaching policy for $2$ billion steps ($36$ hours) and the two-hand reaching policy for 4 billion steps ($60$ hours) on a total of $32,768$ parallel environments. As a result of the robustness introduced by the force perturbations, our model directly transfers to the full humanoid model (with original collision meshes) on the classical MuJoCo simulation, and is also robust to the additional masses introduced by the presence of dexterous hands. We provide the MuJoCo MJX training code as part of our software package.
@@
As mentioned above, we use the reaching policies obtained through MJX training as low-level policies of our hierarchical architecture. For the push and package tasks, we restrict the range of reaching targets to the robot workspace to facilitate exploration, and the high-level policies are trained with either DreamerV3 or TDMPC2. Results are shown in Figure 7, where the hierarchical architecture significantly outperforms the end-to-end baselines on the push task, achieving very high success rates. While the low-level policy has undergone additional pretraining, this can be in principle reused across tasks. We note a less pronounced performance improvement in the more challenging package task. While getting closer to picking up the package, the policy still struggles in lifting it (having never experienced it during training), which limits the overall task performance.
These results confirm that the tasks present in our benchmark present challenges that can be addressed with a more structured approach to the learning problem, and we hope this stimulates further directions for future research.

### v2

@@ Abstract
Humanoid robots hold great promise in assisting humans in diverse environments and tasks, due to their flexibility and adaptability leveraging human-like morphology. However, research in humanoid robots is often bottlenecked by the costly and fragile hardware setups. To accelerate algorithmic research in humanoid robots, we present a high-dimensional, simulated robot learning benchmark, HumanoidBench, featuring a humanoid robot equipped with dexterous hands and a variety of challenging whole-body manipulation and locomotion tasks. Our findings reveal that state-of-the-art reinforcement learning algorithms struggle with most tasks, whereas a hierarchical learning approach achieves superior performance when supported by robust low-level policies, such as walking or reaching. With HumanoidBench, we provide the robotics community with a platform to identify the challenges arising when solving diverse tasks with humanoid robots, facilitating prompt verification of algorithms and ideas. The open-source code is available at https://humanoid-bench.github.io.
@@
TABLE I: Comparison of simulated robot benchmarks. Our humanoid robot benchmark tests a variety of complex, long-horizon task with a large action space.
@@ I Introduction
To accelerate the progress of research for humanoid robots, we present the first-of-its-kind humanoid robot benchmark, HumanoidBench, with a diverse set of locomotion and manipulation tasks, providing an accessible, fast, safe, and inexpensive testbed to robot learning researchers. Our simulated humanoid benchmark demonstrates a variety of challenges in learning for autonomous humanoid robots, such as the intricate control of robots with complex dynamics, sophisticated coordination among various body parts, and long-horizon complex tasks.
HumanoidBench provides (1) a simulation environment comprising a humanoid robot with two dexterous hands, as illustrated in LABEL:fig:teaser; (2) a variety of tasks, spanning locomotion, manipulation, and whole-body control, incorporating humans’ everyday tasks; (3) a standardized benchmark to evaluate the progress of the community on high-dimensional humanoid robot learning and control. In fact, HumanoidBench supports generic controller structures, including both learning and model-based approaches [14, 26]. In this paper, we present extensive benchmarking results of the state-of-the-art reinforcement learning (RL) algorithms, which do not require extensive domain knowledge, and a hierarchical RL approach.
The simulation environment of HumanoidBench uses the MuJoCo [60] physics engine. For the simulated humanoid robot, we mainly opt for a Unitree H1 humanoid robot [^1] https://www.unitree.com/h1, which is relatively affordable and offers accurate simulation models [66], with two dexterous Shadow Hands [^2] https://www.shadowrobot.com/dexterous-hand-series/ attached to its arms. Our environment can easily incorporate any humanoid robots and end effectors; thus, we provide other models, including Unitree G1333https://www.unitree.com/g1, Agility Robotics Digit [^4] https://agilityrobotics.com/robots, the Robotiq 2F-85 gripper, and the Unitree H1 hand.
@@ II Related Work
Complex manipulation tasks, such as block stacking [13], kitchen tasks [17], and table-top manipulation [25, 39, 34], have been introduced but are still limited to a combination of pushing, picking, and placing. On the other hand, the IKEA furniture assembly environment [31], BEHAVIOR [55, 32], and Habitat [57] present diverse long-horizon (mobile) manipulation tasks, with their main focus being on high-level planning by abstracting complex low-level control problems, while FurnitureBench [22] introduces a simulated benchmark for complex long-horizon furniture assembly tasks with sophisticated low-level control.
However, most of these benchmarks use a single-arm manipulation setup with either a parallel gripper or a dexterous hand [9, 49], limiting the types of object interactions and not addressing the challenges of coordinating multiple parts of a body [30], e.g., multiple fingers, arms, and legs. Robosuite [70] includes a handful of bimanual manipulation tasks, while more recently Chen et al. [10] and Zakka et al. [67] have introduced additional benchmarks that require coordinating two floating robot hands, i.e., not attached to any arm base.
While bimanual manipulation is one of the key objectives of humanoid robots, most benchmarks in humanoid research have so far focused on the locomotion challenges [8, 28, 45, 3]. In this regard, such simulations have accelerated research on control algorithms [6, 46, 47, 40], ultimately leading to achieve robust humanoid locomotion in the real world [1, 50, 11].
@@
In contrast to prior robotic simulation benchmarks, HumanoidBench presents a broader set of challenges, featuring high-dimensional action spaces and DoFs, resulting from humanoid robots and dexterous hands, and a variety of long-horizon tasks, which cover a comprehensive set of robotic locomotion and manipulation skills, as summarized in Table I.
Finally, we note how in the literature, tasks that require long-term planning with a high-dimensional action space have been addressed with hierarchical reinforcement learning (HRL), which decouples low-level and high-level planning in a reinforcement learning paradigm [33, 4, 42, 29, 17, 30, 48]. In the context of humanoids, we propose an HRL paradigm to show how a specific set of low-level skills (e.g., standing, walking) facilitates learning of higher-level tasks.
Figure 0: Example egocentric visual (top-left) and whole-body tactile (right) observations when the humanoid interacts with a package in the truck environment. In the right figure, the two cameras on the robot head are highlighted in green, while continuous tactile pressure readings are indicated in shades of red (strong pressure) and yellow (mild pressure). Note that for ease of visualization, we are not showing shear forces and tactile readings on the back of the robot, which are also implemented in our environment.
@@
Figure 1: HumanoidBench manipulation task suite. We devise $15$ benchmarking whole-body manipulation tasks that cover a wide variety of interactions and difficulties. This figure illustrates an initial state for each task (left) and examples of the robot performing such tasks (right).
@@
Figure 2: HumanoidBench locomotion task suite. We devise $12$ benchmarking locomotion tasks that cover a wide variety of interactions and difficulties. This figure illustrates an initial state for each task (left) and examples of the robot performing such tasks (right).
@@ III Simulated Humanoid Robot Environment
In this section, we describe our simulated environment and discuss relevant design choices for the simulated humanoid robot. As illustrated in Figure , we use the Unitree H1 humanoid robotFootnote 1 with two dexterous Shadow HandsFootnote 2 as the primary robotic agent of our benchmark. We simulate this humanoid robot using MuJoCo [60] adapting the Unitree H1 model provided by Unitree [^5] https://github.com/unitreerobotics/unitree_ros and the dexterous Shadow Hand models available through MuJoCo Menagerie. [^6] https://github.com/google-deepmind/mujoco_menagerie
Humanoid Body. We implement Unitree H1Footnote 1, Unitree G1Footnote 3, and Agility Robotics DigitFootnote 4, which are well-known humanoid robots with their model files freely available [66, 1]. Unitree H1 is primarily used in our benchmark as it is a full-size humanoid compared to the smaller Unitree G1, and as we observed faster learning compared to the Agility Robotics Digit, which we ascribe to a simpler mechanical design compared to Digit, which features passive joints actuated through a four-bar linkage.
Dexterous Hands. We use two dexterous Shadow HandsFootnote 2, which also have model files freely availableFootnote 6, and have shown impressive manipulation capabilities both in simulation [67] and in the real world [2]. To make the simulated robot have more human-like morphology, we remove the cumbersome forearms of the dexterous Shadow Hands in HumanoidBench. While this is not currently a realistic model, we anticipate the trend in the industry towards developing slimmer, human-like hands (e.g., Tesla Optimus, Figure 01) so that our design choice aligns better with next-generation humanoid robots. In addition, we also provide models for the Robotiq 2F-85 parallel-jaw gripper and the $13$-DoF Unitree hands available in the Unitree collectionFootnote 5 (see Appendix, Appendix A for more details).
The observation and action spaces, and degrees of freedom of the robot system with or without the dexterous hands are summarized in Table II.
Without hand With $2$ hands
Observation space $51$ $151$
Action space $19$ $61$
DoF (body) $25$ $25$
DoF (two hands) 00 $50$
TABLE II: Humanoid robot specifications with and without hands. Both the humanoid body (including its floating base) and one Shadow Hand present action spaces ($19$ and $21$, respectively) smaller than their DoFs ($25$), making them underactuated systems. In this table, the observation spaces solely comprise generalized positions and velocities of the robots and do not take into account any environment observations. We use quaternions for the robot floating base orientation, which adds an additional position coordinate compared to the velocity components, which match the DoFs. In the appendix, Table III shows an exhaustive overview of all the robot configurations available in HumanoidBench.
@@
  • Proprioceptive robot state (i.e., joint angles and velocities) and task-relevant environment observations (i.e., object poses and velocities).
@@
  • Whole-body tactile sensing using the MuJoCo tactile grid sensor (see Figure ). We design tactile sensing at the hands with high resolution and in other body parts with low resolution, similar to humans, with a total of $448$ taxels spread over the entire body, each providing three-dimensional contact force readings. Similar distributed force readings have been captured on real-world systems both on humanoid bodies [41] and end-effectors [53]. The implementation of such spatially distributed contact sensing required non-trivial mesh adaptations and refinements, which we detail in Appendix, Section B-D.
Although other sensory inputs are available from the environment, to investigate challenges in whole-body control of humanoid robots, we first focus on the state-based environment setup, where proprioceptive robot states and object states are used as the agent’s input in HumanoidBench. In our state-based environment, we maintain the robot observations the same across tasks to minimize domain knowledge, in contrast to tailoring it to the specific tasks [59]. We leave extending our environment to benchmarking multimodal perception capabilities [65, 54] of humanoid robots as future work.
@@
Figure 3: Learning curves of RL algorithms (locomotion). The curves are averaged over three random seeds and the shaded regions represent the standard deviation. Returns are computed by summing the rewards at all timesteps of an episode. The dashed lines qualitatively indicate task success. We run PPO on the walk task but it is not visible in the plot since it only achieves very low returns.
@@ IV HumanoidBench
To this end, we present HumanoidBench, a humanoid benchmark for robot learning and control, which features a high-dimensional action space (up to $61$ different actuators) and enables research in complex whole-body coordination.
We benchmark $27$ tasks, consisting of $12$ locomotion tasks and $15$ distinct manipulation tasks, as illustrated in Figure 2 and Figure 1. A set of locomotion tasks aim to provide interesting but simpler humanoid control scenarios, bypassing intricate dexterous hand control. On the other hand, whole-body manipulation tasks render a comprehensive evaluation of the state-of-the-art algorithms on challenging tasks with unique challenges that require coordination across the entire robot body, ranging from toy examples (e.g., pushing a box on a table) to practical applications (e.g., truck unloading, shelf rearrangement).
In this section, we briefly describe the tasks in our benchmark task suite. Further details about each of the tasks, including task initialization, reward functions, as well as different variations of the tasks, are provided in Appendix, Section B-E.
@@ IV-B Whole-Body Manipulation Tasks
  • cabinet: Open four different types of cabinet doors (e.g., hinge doors, sliding door, drawer).
@@
  • powerlift: Lift a barbell shaped object of a designated mass.
@@ V Benchmarking Results
To identify the challenges in learning with humanoid robots, we benchmark reinforcement learning (RL) algorithms on HumanoidBench, which promises for robots to learn from their own experience. Remarkably, this class of algorithms requires limited domain expertise and does not necessarily rely on expert demonstrations, which are not only expensive but also challenging to collect for humanoid robots. [^7] While we do not benchmark classical model-based control approaches [14, 26] in this work, our environments support actions obtained by using any type of controllers.
@@ V-A Baselines
We evaluate all tasks in our benchmark with four RL methods (DreamerV3, TD-MPC2, SAC, PPO). Please refer to Appendix, Appendix C for implementation details.
@@ V-B Results
We report benchmarking results in Figure 3 and Figure 4, where we ran each of the algorithms for approximately $48$ hours, resulting in the visible differences in environment steps (e.g., $2$M steps for TD-MPC2, $10$M steps for DreamerV3). We only run PPO on a subset of tasks (walk, kitchen, door, package), given its inferior performance without massive parallelization. Each of the environments is evaluated with a combination of dense rewards and sparse subtask completion rewards, and for each of these we provide qualitative measures of task success (see dashed lines in Figure 3 and Figure 4). A detailed description of the reward functions used for each environment is available in Appendix, Section B-E.
@@
This poor performance is mainly attributed to the high-dimensionality of the state and action spaces of our humanoid robot agent with dexterous hands. Although the hands of the humanoid robot are barely used for most locomotion tasks, the RL algorithms fail to ignore this information, which makes policy learning challenging. In addition, these high-dimensional state and action spaces result in a much larger exploration space, which makes exploration slow or infeasible with simple maximum entropy approaches. This implies the need for incorporating behavioral priors or commonsense knowledge about the world that can ease the exploration problem, when it comes to learning on more complex agents, like humanoid robots. We investigate this further in Section V-C.
This problem becomes even more severe in manipulation tasks, resulting in particularly low rewards in all such tasks. Before learning any manipulation skills, an agent must learn locomotion skills to balance and move towards an object or the world to interact. All the policies barely learn to stabilize using the dense reward, but struggle to learn any complex manipulation skills.
@@
V-C With Hands vs. Alternative Configurations
With Hands vs. Without Hands. As discussed in Section V-B, controlling humanoid robots with dexterous hands is challenging due to their high degrees of freedom and complex dynamics. Thus, we investigate the difficulty of RL training with a large action space (i.e., additional $42$ dimensions with two dexterous Shadow Hands) on walk that does not necessarily require to control dexterous hands. The results in Figure 5 show that the presence of hands, with their additional joints and actuators, leads to a large decrease in performance compared to training the same task without the dexterous hands (see differences in observation and action space in Table II).
Reduced Action Space. To verify whether such difficulties stem from the dimensionality of the action space, we benchmark our full robot model, but fix the actuation of the hands ($42$D), which we set to zero. In this way, the action dimensionality is reduced from $61$D in the original model to $19$D. Note that the observations and masses induced by the presence of the hands are retained (i.e., observation space remains $151$D). Figure 5 shows that the RL algorithms learn significantly faster in the reduced action space setup than the ones trained with the full action space. This confirms that most of the performance drop is indeed due to the increased action dimensionality.
We observe similar trends in the more complex manipulation task, push, which presents substantially different dynamics in the task approach (e.g., pushing with and without hands).
@@
(a) Hierarchical learning pipeline
(b) Low-level policy pretraining
(c) High-level policy training
Figure 6: Our hierarchical RL pipeline (a). (b) A robust low-level reaching policy is pretrained using PPO in a MuJoCo MJX-based reaching environment, as shown in the top snapshot in (a). (c) The high-level policy then leverages the pretrained reaching policy to move to a desired position and learns to solve a downstream task, shown in the bottom snapshot in (a). Note that the reaching policy weights are frozen during the high-level policy training.
@@
As shown in the previous subsection, flat, end-to-end RL approaches fail to learn most of the tasks in HumanoidBench. Many of such tasks require long-horizon planning and necessitate acquiring a diverse set of skills (e.g., balancing, walking, reaching). These tasks can be addressed by hierarchical RL, which introduces additional structure into the learning problem. In HRL, one or multiple low-level skill policies are provided to a high-level planning policy that outputs setpoints for the lower-level policies. In practice, such setpoints comprise the action space of the high-level policy. This framework is very general, and there are no constraints on how to obtain both low-level and high-level policies. However, here we focus on training both of these through reinforcement learning [56].
@@
Hierarchical RL Implementation. We implement a hierarchical RL approach on two manipulation tasks, namely, the push and package tasks. As a low-level skill, push uses a one-hand reaching policy, which allows the robot to reach a 3D point in space with its left hand, while package uses a two-hand reaching policy, where both hands are commanded to reach different 3D targets. Figure 6 illustrates the overview of our hierarchical RL implementation.
@@
Low-level Reaching Policy Pretraining. We treat the low-level reaching policy as a pretrained frozen block that can be reused across tasks. Since this policy does not improve during training of the high-level policy, it needs to be very robust to cope with the continually shifting reaching targets that the high-level policy sets during exploration. However, the results in the previous section show that even a one-hand reaching task is hard to learn.
On the other hand, while our experiments above confirm that PPO exhibits poor sample efficiency compared to the off-policy algorithms, it is worth noting that PPO has achieved significant success in robotic locomotion by exploiting large-scale parallelization of environments on GPUs [36]. To achieve robust reaching policies, we exploit hardware acceleration by pretraining the low-level reaching policies in the recently released MuJoCo MJX [^8] https://mujoco.readthedocs.io/en/stable/mjx.html, which enables training PPO on thousands of parallel environments.
@@
For low-level reaching policy training, we employ a simplified H1 model that only considers collisions between feet and ground in the MuJoCo MJX environments, as in our experience the advantages stemming from parallelization are largely reduced when considering all numerous humanoid geometries (hindering training of the more complex benchmark tasks via MJX). We also remove the hands from the model to further increase training efficiency. The simplified reaching task environments for pretraining reset the target once reached. To achieve robust low-level reaching policies, we apply force perturbations at each of the links during training. We train the one-hand reaching policy for $2$ billion steps ($36$ hours) and the two-hand reaching policy for $4$ billion steps ($60$ hours) on $32{,}768$ parallel environments. The pretrained reaching policies successfully transfer to the original (non-simplified, simulated in classical MuJoCo) humanoid environments.
High-level Policy Training. Then, we use the pretrained reaching policies (frozen) as low-level policies and only train a high-level policy using either DreamerV3 and TD-MPC2 on the push and package tasks. To facilitate exploration, we restrict the range of reaching targets to the robot workspace.
Hierarchical RL Results. In Figure 7, our hierarchical architecture significantly outperforms the flat, end-to-end baselines on the push task, achieving very high success rates with DreamerV3. While the low-level policy has undergone additional pretraining, this can be in principle reused across tasks. On the other hand, we note a less pronounced performance improvement in the more challenging package task. While getting closer to picking up the package with our hierarchical approach, the policy struggles in lifting it (having never experienced it during training).
@@
These results confirm that the tasks in our benchmark present challenges that can be addressed with a more structured approach to the learning problem, and we hope this stimulates further directions for future research.